OpenStudy (anonymous):

Write the equation of the circle with center (-6, -4) and containing the point (-2, -1).

OpenStudy (amistre64):

given a center point (a,b) (x-a)^2+(y-b)^2 = r^2 such that r is the distance between the center and the given point

OpenStudy (anonymous):

What do I do with the contained point (-2,-1)?

OpenStudy (amistre64):

that helps to define the radius of your circle the distance from the center to that point ....

OpenStudy (amistre64):

do you recall how to find the distance between 2 points?

OpenStudy (anonymous):

yes by using the distance formula which is\[\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

so i take both the contained point and the center and plug it in ti that equation?

OpenStudy (amistre64):

well, seeing that we need to define the distance between them; and that is the formula for finding the distance between 2 points ... im going with yes :)

OpenStudy (anonymous):

Wonderful

OpenStudy (amistre64):

the formula is nice but i tend to place things in the wrong spots .... so i just step thru it subtract the points (-6, -4) -(-2, -1) -------- -4, -3 squre that parts: 16, 9 add them: 16+9 = 25 and sqrt: sqrt(25) = 5

OpenStudy (anonymous):

So after substituting the two points into the distance formula i get a radius of 5.385164807

OpenStudy (amistre64):

lets go with 5

OpenStudy (anonymous):

which then squared equals 29

OpenStudy (anonymous):

Or we could do 5...

OpenStudy (amistre64):

putting it all together we should get \[(x+6)^2+(y+4)^2=25\]

OpenStudy (anonymous):

That is what I got!

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

Thanks for your help my friend.

OpenStudy (amistre64):

youre welcome